[Brahms-l] Shift report 20050216 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Feb 16 2005 - 00:02:42 EST
Shifter leader was hy
Shifters were 


1. Machine development for almost the whole shift, at the end of the shift, they tried to get beam for physics, but they lost it after few mimutes. 
2. run 13200 and run 13201 for pedestal run
3. D1 problem, we can not control it from CH. see the highlighted part in the e-log for information in detail.  
When I was doing the shift check list, HV of most channels in TRMRS, TOFW and CALFB are OFF actually. Called Hiro, he fixed it. The reason is the following: Lost communication to MRS 1440 HV main frame. 





_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Feb 16 00:03:20 2005

This archive was generated by hypermail 2.1.8 : Wed Feb 16 2005 - 00:03:35 EST